# A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection

### v3

@@ II.2.3 Formulations
We now derive the DE and DCR for the $(s+1)$-th level.
For the DE calculation, we assume a non-vacuum input and calculate the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
@@
Combining these, the DE of $\text{ESPD}_{s+1}$ is:
@@
For the DCR calculation, we assume a vacuum input. The probability of a false positive is:
$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@
II.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, employing the approximations in Eq.(1) to (3) and the assumptions from Section II.2.2, we can derive simplified bounds to elucidate the system’s behavior.
Consider the DCR evolution. Eq. (7) can be reformulated as:
$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@ II.3 Analytical Approximations and Stability Analysis
Then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
@@
$\frac{d}{dx}f(x)$ (11)
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
Applying the upper bound from Eq. (2), substituting Eq. (3) to Eq. (9), and assuming $k-1\geq nx$, we obtain:
@@
which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate $n_{s+1},k_{s+1}$, we assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
@@
which is independent of $d_{s}$. Therefore, solving the inequality
@@
determines the condition under which the DE is enhanced in the subsequent level. If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
II.4 Performance Simulations
@@
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations for the ESPD paradigm, indicating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs. The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section III.1 for parameter discussions). We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section III.1).
Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2 and 3. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
II.5 Application on QKD
To illustrate the practical impact of the ESPD scheme, we consider its application in QKD. For a QKD protocol having a secure threshold $e_{th}$ and total error rate $e_{C}$ not accounting for SPDs, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
@@ II.5 Application on QKD
Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude (even compared to SNSPDs), it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor. Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96. Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section II.4.
@@
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency. This allows significant QKD performance improvements without additional efficiency costs. Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
III Discussion
In the absence of experimental implementation, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations and directions for future work.
III.1 Technological Feasibility and Performance under Degraded Parameters
III.1.1 State Preparations and Measurements
The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates. Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
III.1.2 Controlled Gates
@@
Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section II.4.
@@
A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section II.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.4 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
@@
Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42]. However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6). All other settings remain consistent with those described in Section II.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
III.1.3 Choice of initialized SPDs
@@
It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation. As supported in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations suggest that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
@@
In summary, the ESPD scheme only requires a normal, room-temperature SPD for initialization, which is routinely achievable. For instance, even an SPD developed two decades ago [28] is sufficient.
III.1.4 Auxiliary Sources

### v4

@@ 2.2.3 Formulations
The DE and DCR for the $(s+1)$-th level are derived as follows.
For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
@@
By combining the above contributions, the DE of $\text{ESPD}_{s+1}$ is given by:
@@
For the DCR calculation, a vacuum input is assumed. The probability of a false positive is given by:
$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
@@
2.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, by employing the approximations in Eq. (1) to (3) together with the assumptions in Section 2.2.2, simplified bounds can be derived to elucidate the system behavior.
The evolution of the DCR is considered first. Eq. (7) can be reformulated as:
$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (9)
@@ 2.3 Analytical Approximations and Stability Analysis
for $a\geq 0$, then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
@@
$\frac{d}{dx}f(x)=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$ (11)
$+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, the following bound is obtained:
@@
which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate values of $n_{s+1}$ and $k_{s+1}$, it is assumed that $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
@@
which is independent of $d_{s}$. Therefore, the condition for DE enhancement at the subsequent level is determined by solving the inequality
@@
If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
2.4 Numerical Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, numerical simulations of the ESPD paradigm are presented, indicating that substantial performance gains can be achieved even with moderate initial hardware specifications. For example, a three-level ESPD can upgrade a semiconductor SPD to effective performance metrics with DE $93\%$ and DCR $10^{-9}$, comparable to those of state-of-the-art SNSPDs without reliance on superconducting materials.
Simulation Parameters: The performance of the ESPD scheme is evaluated using two distinct baseline SPDs. The first represents an older one, characterized by parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section 3.1 for parameter discussions). Supplementary simulations with degraded parameter settings are also performed to evaluate robustness in Section 3.1.
Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
@@ 2.4 Numerical Simulations
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
(d) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
Figure 2: Performance of ESPD initialized with SPDs from [28] ((a), (b)) and [29] ((c), (d)), respectively.
2.5 Implications for QKD
As a representative application scenario, the implications of the ESPD framework for QKD are analyzed. For a QKD protocol characterized by a secure error threshold $e_{th}$ and a total error rate $e_{C}$ excluding detector contributions, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by[27]:
@@ 2.5 Implications for QKD
Since the ESPD paradigm can substantially suppress the effective DCR $d$, the minimal tolerable channel transmission rate $\gamma$ is proportionately reduced. Furthermore, the simultaneous enhancement of the DE $\eta$ not only reduces $\gamma$ but also directly improves the achievable key rate at a fixed channel transmission rate.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD framework achieves this improvement without introducing additional sifting procedures that reduce the effective sifting efficiency. Here, QKD performance can be enhanced through the simultaneous suppression of the effective DCR and enhancement of the effective DE, without the need for auxiliary sifting operations. In this sense, the ESPD paradigm provides a more efficient system-level approach for improving QKD performance under practical constraints.
3 Discussion
In the absence of an experimental demonstration, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations as well as directions for future work.
3.1 Technological Feasibility and Performance under Degraded Parameters
3.1.1 State Preparations and Measurements
The flexibility of the ESPD architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be achieved at extremely low levels. Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
3.1.2 Controlled Gates
Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96. Simulations are performed for $(n,k)=(5,2)$ at each level, and all other settings remain consistent with those described in Section 2.4.
@@ 3.1.2 Controlled Gates
Figure 3: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, does not decrease with an increasing number of cascaded modules or ESPD levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.3 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryonic components[39, 40, 41, 42]. However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), with visual representation in Fig.4. All other settings remain consistent with those described in Section 2.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with existing room-temperature quantum technologies.
@@
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.40$.
(d) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$, with $P=0.40$.
@@
Figure 4: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$ ((a), (b)) and $P=0.40$ ((c), (d)), respectively.
3.1.3 Choice of initialized SPDs
@@
Notably, the requirements imposed on the SPD used for initialization are minimal, which further lowers the practical barrier to implementation. As evidenced by the comparison between Tables 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the initialized SPD, provided that it lies within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, numerical simulations suggest that both SPDs with low DE and low DCR[29], as well as those with moderate DE and relatively high DCR[28], can serve as viable initializations.
@@
In summary, the ESPD scheme requires only a conventional room-temperature SPD for initialization, which is routinely achievable with existing technology. For instance, the SPD developed two decades ago [28] is sufficient.
3.1.4 Auxiliary Sources